use the method of substitution to solve the system of linear equations.
2x-y=4
9x-4y=19
The solution is ______
Please help me solve this.

Answer by jim_thompson5910(35256)

2x-y=4

-y=4-2x

-y = -2x + 4

y = 2x - 4


9x - 4y = 19

9x - 4(2x - 4) = 19

9x - 8x + 16 = 19

x + 16 = 19

x = 19 - 16

x = 3

Now that you have the solution for x, use this to find the solution for y.